Solve equation, check answer
root 3√(2x-3) + 1 = 4

Answer by nerdybill(7384)

root 3√(2x-3) + 1 = 4
root 3√(2x-3) = 3
2x-3 = 3^3
2x-3 = 27
2x = 30
x = 15